# When Unsupervised Domain Adaptation meets One-class Anomaly Detection: Addressing the Two-fold Unsupervised Curse by Leveraging Anomaly Scarcity

### v1

@@ 2 Related Works: Anomaly detection under domain shift
$\min_{\zeta}\,\mathbb{E}_{(\mathbf{X}_{i},y_{i})\sim\mathcal{D}^{n}}\left[\mathcal{L}\left(\zeta(\mathbf{X}_{i}),y_{i}=0\right)\right],$ (1)
where $\mathcal{L}$ is a loss enforcing feature compactness as in DSVDD [28] or a reconstruction loss typically used in autoencoders-based methods [7, 35]. Although achieving impressive performance on standard benchmarks, the majority of AD methods [11, 29, 13, 28, 7, 35, 33] overlook the domain gap problem where training and testing data denoted as $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$, respectively, follow different distributions due to uncontrolled variations in the acquisition setting [4, 5]. This domain shift induces, therefore, a significant drop in performance. To solve this issue, a handful of domain generalization methods for UAD has been proposed recently [5, 4]. In [5], multiple source domains are considered for learning domain-invariant features, thereby assuming the availability of diverse large-scale datasets which is not always guaranteed. To avoid relying on multiple domains during training, a self-supervised strategy is adopted in [4]. Nevertheless, the success of this approach heavily depends on the similarity between the augmented data and target samples. As a result, it necessitates tailoring augmentation techniques to unseen target datasets, if at all possible. Given its effectiveness, domain adaptation has also been explored to address the domain shift problem in AD [6, 18, 16, 19, 20]. Those techniques usually adopt a few-shot adaptation paradigm by having access to a limited number of annotated target samples. While these methods offer innovative solutions for aligning source and target normal data, they still rely on costly annotations [13] and are exposed to overfitting risks [32]. This emphasizes the need for a fully unsupervised domain adaptation for UAD. However, solving this challenge remains challenging given the doubly unsupervised nature of the problem resulting from both UAD and UDA, which is further described in the next section.
@@ 5 Proposed Solution to the Two-fold Unsupervised Curse
Our approach has two branches. The upper branch depicts a trainable backbone $f$ that learns from both source and target domain data. The source features are optimized using a Deep Support Vector Data Description (DSVDD) objective [28]. The lower branch focuses on visual feature extraction from the unlabeled target domain through a frozen CLIP visual encoder [27], defined as the $\psi$ feature extractor. Clustering is applied to these visual features to estimate the dominant cluster $\mathcal{C}$. Samples identified within $\mathcal{C}$ in the $\psi$ visual encoder’s representation space are then selected within the space of the feature extractor $f$ and then aligned with the normal source features.
Specifically, given source and target image datasets $\mathcal{D}^{s,n}$ and $\mathcal{D}^{t}$, we apply DSVDD on the source data, enforcing feature compactness by minimizing the radius of a hypersphere to encapsulate the normal source representations. This is done by solving the following optimization problem,
@@
where $\boldsymbol{\mu}^{s,n}$ is the mean of the source features. For clustering, we use a $K$-means algorithm. Note that $\psi$ can be $f$ itself or any frozen visual encoder such as CLIP [27] or DINO-v2 [24].
The dominant cluster is identified as,
@@
$\ell_{i,j}=-\log\frac{\exp\left(\frac{\text{sim}\left(f(\mathbf{X}_{i}^{s}),f(\mathbf{X}_{j}^{t})\right)}{\tau}\right)}{\sum_{p=1}^{N_{t}}\mathbbm{1}_{[\mathbf{X}_{p}^{t}\notin\mathcal{\tilde{D}}^{t,n}]}\exp\left(\frac{\text{sim}\left(f(\mathbf{X}_{i}^{s}),f(\mathbf{X}_{p}^{t})\right)}{\tau}\right)},$ (8)
@@
At inference, note that the visual encoder $\psi$ is discarded and that only the feature extractor $f$ is used for estimating whether the input data falls inside or outside the hypersphere estimated by the DSVDD model.
@@ 6 Experimental Results
CAD $\rightarrow$ Real
Aeroplane 67.71 36.80 81.56 81.55 84.86
@@
Feature Extractor w/ Adaptation w/ Clustering w/ CLIP AUC (%)
Resnet18 56.80$\pm$10.90
56.08$\pm$7.22
63.43$\pm$8.39
72.47$\pm$11.04
Resnet50 62.84$\pm$8.55
@@
Normal class Pretrained Visual encoders Source only (finetuned) Few-shot Adaptation Unsup. Adapt. Oracle (Supervised)
CLIP-R50 CLIP-ViT-B32 BiOST MsRA Ours Ours
CAD $\rightarrow$ Real
Aeroplane 33.97 74.97 67.71 36.80 81.56 81.55 84.86 90.91
Bicycle 53.52 90.28 65.12 59.20 68.45 74.58 81.45 81.73
Bus 46.40 42.27 66.01 47.90 68.12 72.26 82.17 72.16
Car 67.84 64.16 78.65 53.80 69.44 82.78 62.76 68.42
Horse 44.80 75.48 67.24 58.00 68.77 80.17 83.52 88.70
Knife 67.76 95.28 62.43 54.10 70.39 71.52 68.82 78.90
Motorcycle 61.22 82.25 69.45 58.10 65.64 80.16 91.15 83.46
Person 56.50 56.26 42.11 58.70 59.18 51.24 69.68 85.19
Plant 62.06 89.65 57.77 42.10 65.81 71.46 70.58 82.63
Skateboard 67.97 91.52 60.70 41.60 61.30 63.17 83.71 83.73
Train 53.61 57.74 54.75 52.40 69.73 60.62 69.98 85.11
Truck 44.82 45.08 62.08 43.10 59.05 73.67 57.84 78.91
Avg.$\pm$±std 55.04$\pm$10.47 72.08$\pm$17.83 62.84$\pm$8.55 50.48$\pm$7.55 67.28$\pm$5.79 71.93$\pm$9.08 75.54$\pm$9.80 81.65$\pm$6.11
Table 4: Performance in AUC (%) of our Resnet50-based one-class anomaly detector on the target domain of the VisDA dataset [25] using various adaptation paradigms (from zero-shot to supervised).
Method Ph. $\rightarrow$ Art Ph. $\rightarrow$ Cartoon Ph. $\rightarrow$ Sketch
Source only 64.06 64.08 57.35
GNL [4] 65.62 67.96 62.39
Ours 67.20 75.35 74.04
Table 5: AUC (%) performance comparison with domain generalization for anomaly detection trained on the PACS [17] where Photo (Ph.) is used as the source domain.
@@
Dataset Source only w/ Adaptation
Kmeans GMM Meanshift
VisDA 62.84$\pm$08.55 75.54$\pm$10.23 72.24$\pm$08.81 74.14$\pm$07.44
[…]
Table 7: Ablation on different clustering techniques. GMM and K-means use $10$ components for VisDA and $2$ for other datasets.
After presenting the experimental setting in Section 6.1, we compare experimentally our method against prior works in Section 6.2. Finally, we conduct additional experiments to assess the contributions of individual components within the proposed method in Section 6.3.
6.1 Experimental Setting
This section describes the datasets and the baselines used in our experiments. The evaluation metrics and the implementation details are provided in supplementary materials. In all the experiments, the best and second performances are highlighted in bold and underlined, respectively.
Datasets.
We evaluate our approach on four standard UDA benchmark datasets, Office-Home [30], Office31 [30],VisDA [25], and PACS [17]. For the AD task, we adopt a standard one-vs-all protocol, where a single class is considered normal and the remaining are anomalies. We adopt the experimental protocol of previous DA works [19, 38, 6] to allow for a fair comparison –that is, we show results on ten classes from the ClipArt and Product domains for Office-Home, ten classes from Webcam and Amazon for Office31, and twelve classes from CAD to Real from VisDA. On PACS [17], like [4], we consider the Photo domain as source and the remaining domains as targets.
Baselines.
As no other works on UDA for semantic image UAD were previously introduced, we compare our method with several few-shot SoA approaches. More specifically, we consider BiOST [6] which is a one-shot approach and TSA [18], ILDR [16], IRAD [38], and MsRA [19] that are few-shot methods. Furthermore, we introduce a few-shot variant (Ours-Few-shot) of our approach, which augments the target domain with normal and pseudo-anomalous samples similar to [9]. This augmentation yields semantically positive and negative pairs [9] useful for the contrastive alignment strategy described in Section 5. Additional details are provided in the supplementary materials.
6.2 Comparison against State-of-the-art.
Our method outperforms previous SoA on the three considered benchmarks used in our evaluation, as seen in Table 1 and Table 2. More specifically, our fully unsupervised UDA-UAD importantly improves upon previous few-shot SoA on C $\rightarrow$ P and W $\rightarrow$ A of the Office-Home [36] and Office31 [30] datasets. In addition, we observe an improvement of over $10\%$ in the VisDA dataset [25] with the fully unsupervised methodology over previous few-shot approaches despite being challenged by the two-fold unsupervised curse. These results highlight the relevance of the proposed method, even in the presence of a large domain gap, as in the case of CAD images and real-world photos.
Furthermore, we compare the performance of our model to two pretrained CLIP visual encoders [27], namely CLIP-R50 and CLIP-ViT-B32 in  Table 4. While the larger CLIP-ViT-B32 architecture achieves an average AUC of 72.08%, our few-shot variant (71.93%) and unsupervised method (75.54%) still outperform it on the VisDA dataset [25]. In contrast, the smaller CLIP-R50 model gives significantly lower performance, with an average AUC of 55.04%. These results demonstrate that despite their strong performance, pretrained visual encoders are not specifically tailored for the domain adaptation task; thus, they remain vulnerable to domain shift. Therefore, training domain adaptation-specific models is still necessary to effectively bridge the gap between two given domains.
6.3 Additional Experiments
Unless stated otherwise, all the following experiments are performed on VisDA [25]. We also report additional results and implementation details in the supplementary materials.
Few-shot versus unsupervised paradigms.
The results presented in Table 4 compare pretrained visual encoders and source-only detectors with different adaptation paradigms, i.e., few-shot, unsupervised, and supervised (oracle). The source-only finetuned model improves slightly over the pretrained CLIP-R50 visual encoder [27] but still has lower performance than the adaptation approaches, achieving an average AUC of 62.84%. Among the few-shot methods, our few-shot variant outperforms BiOST [6] and MsRA [19], achieving the highest AUC of 71.93%, which is comparable to the performance of a pretrained CLIP-ViT-B32 visual encoder. However, our unsupervised adaptation method surpasses all these models, with an average AUC of 75.54%, demonstrating its ability to effectively mitigate domain gaps without relying on labeled target data. This can be explained by the fact that after clustering, our model has access to more representative normal target data than few-shot models, hence better generalizing to the target normal class. On the other hand, the oracle-supervised model, with access to full target labels, achieves the highest performance (81.65%). The relatively small gap between our unsupervised method and oracle confirms the relevance of the proposed unsupervised framework.
Ablation on the framework components.
Table 3 provides the results obtained when each component, namely the adaptation loss, the dominant cluster identification through clustering, the use of an auxiliary visual encoder $\psi(\mathcal{D}^{t})$ or the trainable features $f(\mathcal{D}^{t})$. The results show that without adaptation, a model trained only on source data generalizes poorly to the target domain with only 56.80% and 62.84% for ResNet18 and ResNet50, respectively. Direct adaptation of the source and the unlabeled target without clustering leads to inconsistent results, indicating low generalization capabilities to the target domain. Introducing clustering results in a significant performance boost. This can be seen when clustering is applied to the original representations of the feature extractor, as the performance of ResNet18 improves by +6.63% and ResNet50 by +5.63%, highlighting the importance of identifying the dominant cluster prior to alignment. Note that our method using clustering and contrastive alignment strategy still outperforms the best few-shot baseline MsRA [19]. Finally, the best results are achieved when all components are combined. This configuration boosts the AUC to 72.47% for ResNet18 and 75.54% for ResNet50. The substantial performance gains can be attributed to the CLIP’s rich visual features, which, together with clustering and alignment, help achieve a more robust anomaly detector capable of effectively handling domain shift.
Figure 4: Effect of varying the number of components in K-Means.
Clustering methods.
We compare different clustering techniques on all three UDA benchmarks in Table 7. The first observation we make is that any type of clustering contributes to improving the performance. K-means and GMM have comparable results, without one clearly and consistently outperforming the other across datasets and adaptation directions. Meanshift clustering offers a performance increase compared to source-only models. However, its performance remains lower than that of the other clustering methods. In our experiments, we choose K-means clustering as it offers the best performance improvement with fewer parameters and simpler optimization than GMM. We further investigate the optimal number of K-Means components, as shown in Figure 4. The figure indicates that using 8 to 10 components yields the highest performance, with an AUC of approximately 75-76%. Decreasing the number of components would gradually degrade the performance. This suggests that a lower number of clusters may not capture the characteristics of the majority class, leading to inaccurate clustering and thus negatively impacting the generalization of the anomaly detection model across domains.
Alignment strategies.
@@ 6.3 Additional Experiments
Table 5 compares the results of the GNL [4] with proposed UDA method on the PACS dataset [17], with Photo as the source and Art, Cartoon, and Sketch as the target domains. It can be seen that our UDA approach consistently outperforms the GNL [4] method, particularly on Cartoon and Sketch domains. This suggests the suitability of UDA for semantic UAD as as it exposes the model to the target domain even if unlabeled.

### v2

@@ 2 Related Works: Anomaly detection under domain shift
$\min_{\zeta}\,\mathbb{E}_{(\mathbf{X}_{i},y_{i})\sim\mathcal{D}^{n}}\left[\mathcal{L}\left(\zeta(\mathbf{X}_{i}),y_{i}=0\right)\right],\vspace{-1mm}$ (1)
where $\mathcal{L}$ is a loss enforcing feature compactness as in DSVDD [33] or a reconstruction loss typically used in autoencoders-based methods [8, 40]. Although achieving impressive performance on standard benchmarks, the majority of AD methods [13, 34, 16, 33, 8, 40, 38] overlook the domain gap problem where training and testing data denoted as $\mathcal{D}^{s}$ and $\mathcal{D}^{t}$, respectively, follow different distributions due to uncontrolled variations in the acquisition setting [4, 5]. This domain shift induces, therefore, a significant drop in performance. To solve this issue, a handful of Domain Generalization (DG) methods for UAD has been proposed recently [6, 5, 4]. Cohen et al. [6] propose a domain-disentanglement approach that removes predefined nuisance attributes (e.g., pose, lighting) from the source features using contrastive loss, preventing these factors from interfering with the anomaly task, improving the performance on unseen domains. However, without an actual target set, this method requires defining and labeling nuisance factors within the source dataset, which is challenging, as mentioned in their paper. In [5], multiple source domains are considered for learning domain-invariant features, thereby assuming the availability of diverse large-scale datasets, which is not always guaranteed. To avoid relying on multiple domains during training, a self-supervised strategy is adopted in [4]. Nevertheless, the success of this approach heavily depends on the similarity between the augmented data and target samples. As a result, it necessitates tailoring augmentation techniques to unseen target datasets, if at all possible. Given its effectiveness, domain adaptation has also been explored to address the domain shift problem in AD [7, 21, 19, 22, 23]. Those techniques usually adopt a few-shot adaptation paradigm by having access to a limited number of annotated target samples. While these methods offer innovative solutions for aligning source and target normal data, they still rely on costly annotations [16] and are exposed to overfitting risks [37]. This emphasizes the need for a fully unsupervised domain adaptation for UAD. However, solving this challenge remains challenging given the doubly unsupervised nature of the problem resulting from both UAD and UDA, which is further described in the next section.
@@ 5 Proposed Solution to the Two-fold Unsupervised Curse
Our approach has two branches. The upper branch depicts a trainable backbone $f$ that learns from both source and target domain data. The source features are optimized using a Deep Support Vector Data Description (DSVDD) objective [33]. The lower branch focuses on visual feature extraction from the unlabeled target domain, through a frozen CLIP visual encoder [30], defined as the $\psi$ feature extractor. Clustering is applied to these visual features to estimate the dominant cluster $\mathcal{C}$. Samples identified within $\mathcal{C}$ in the $\psi$ visual encoder’s representation space are then selected within the space of the feature extractor $f$ and then aligned with the normal source features. An algorithm is provided in Section A of the Supplementary.
Training. Specifically, given source and target image datasets $\mathcal{D}^{s,n}$ and $\mathcal{D}^{t}$, we apply DSVDD on the source data, enforcing feature compactness by minimizing the radius of a hypersphere to encapsulate the normal source representations. This is done by solving the following optimization problem,
@@
where $\boldsymbol{\mu}^{s,n}$ is the mean of the source features. For clustering, we use a $K$-means algorithm. Note that $\psi$ can be $f$ itself or any frozen visual encoder such as CLIP [30] or DINO-v2 [27]. The dominant cluster is identified as,
@@
$\ell_{i,j}=-\log\frac{\exp\left(\frac{1}{\tau}\cdot\text{sim}\left(f(\mathbf{X}_{i}^{s}),f(\mathbf{X}_{j}^{t})\right)\right)}{\sum_{p=1}^{N_{t}}\mathbbm{1}_{[\mathbf{X}_{p}^{t}\notin\mathcal{\tilde{D}}^{t,n}]}\exp\left(\frac{1}{\tau}\cdot\text{sim}\left(f(\mathbf{X}_{i}^{s}),f(\mathbf{X}_{p}^{t})\right)\right)}$ (8)
@@
Inference. Note that the visual encoder $\psi$ is discarded at inference and that only the feature extractor $f$ is used to determine whether the input data is anomalous by calculating whether it falls inside or outside the hypersphere estimated by the DSVDD model. Our method’s algorithm is given in Section A of the Supplementary.
@@ 6 Experimental Results
Synthetic (CAD) $\rightarrow$ Real
Aero. 67.71 36.80 81.56 81.55 84.86
@@
6.1 Experimental Setting
This section describes the datasets and the baselines used and the implementation details of our experiments. We report the performance using Area Under the ROC Curve (AUC) using bold and underline for the best and second performances, respectively. The supplementary materials provide further details on each subsection.
Datasets.
We evaluate our approach on four standard UDA benchmark datasets, Office-Home [35], Office31 [35],VisDA [28], and PACS [20]. The types of domain shift of each dataset is described in Section A of the Supplementary. For the AD task, we adopt a standard one-vs-all protocol, where a single class is considered normal and the remaining are anomalies. We adopt the experimental protocol of previous DA works [22, 43, 7] to allow for a fair comparison –that is, we show results on ten classes from the ClipArt and Product domains for Office-Home, ten classes from Webcam and Amazon for Office31, and twelve classes from the domains of Computer Aided Designs (CAD) (synthetic objects) and real object photos of VisDA. On PACS, like [4], we consider the Photo domain as source and the remaining three domains as targets.
Baselines.
As no other works on UDA for semantic image UAD were previously introduced, we compare our method with several few-shot SoA approaches. More specifically, we consider BiOST [7] which is a one-shot approach and TSA [21], ILDR [19], IRAD [43], and MsRA [22] that are few-shot methods. Furthermore, we introduce a few-shot variant (Ours-Few-shot) of our approach, which augments the target domain with normal and pseudo-anomalous samples similar to [11]. This augmentation yields semantically positive and negative pairs [11], which are useful for the contrastive alignment strategy described in Section 5. Further details are given in Section A of the Supplementary.
Implementation details.
In all experiments, the source set has only normal data, while the unlabeled target set includes mostly normals with 10% randomly sampled anomalies. Training uses SGD with a cosine-annealing scheduler, learning rate of $10^{-3}$, weight decay of $5\times 10^{-7}$, batch size 256 and $\lambda_{1}$ and $\lambda_{2}$ are set to 1. CLIP-ViT-B32 is the frozen visual encoder $\psi$ for feature clustering. To align with the setting of the baselines [22, 43, 7, 21], ResNet50 is the trainable backbone $f$, initialized on ImageNet [9]. K-means clustering [14] uses 2, 10, and 5 components for Office, VisDA, and PACS, respectively. Like the baselines, the few-shot experiments use 10 (Office, PACS) and 100 shots (VisDA) labeled as normal, respectively.
6.2 Comparison against State-of-the-art.
Our method outperforms previous SoA on all benchmarks of our evaluation, as shown in Table 1 and Table 2. More specifically, our fully unsupervised UDA-UAD importantly improves upon previous few-shot SoA on C $\rightarrow$ P and W $\rightarrow$ A of the Office-Home [41] and Office31 [35] datasets. In addition, we observe an improvement of over $10\%$ in the VisDA dataset [28] with the fully unsupervised methodology over previous few-shot approaches despite being challenged by the two-fold unsupervised curse. These results highlight the relevance of the proposed method, even in the presence of a large domain gap, as in the case of synthetic CAD images and real-world photos.
Furthermore, we compare the performance of our model to two pretrained visual encoders [30], namely ResNet50 and CLIP-ViT-B32 in Table 3. While the CLIP-ViT-B32 architecture achieves an average AUC of 72.08%, our unsupervised method (75.54%) still outperforms it on the VisDA dataset [28]. In contrast, the ResNet50 shows a significantly lower performance, with an average AUC of only 53.45%. These results demonstrate that despite their strong performance, pretrained visual encoders are not specifically tailored for the domain adaptation task; thus, they remain vulnerable to domain shift. Therefore, training domain adaptation-specific models is still necessary to effectively bridge the gap between two given domains.
6.3 Additional Experiments
Unless stated otherwise, all the following experiments are performed on VisDA [28]. Additional results on different components of our method are given in the Supplementary.
Anomaly scarcity assumption.
To evaluate the impact of anomaly scarcity, we vary the anomaly ratio in the unlabeled target set from 10% to 90% and report our method’s performance alongside clustering accuracy in Figure 4 for the Aeroplane class from VisDA. The results indicate a strong correlation between AUC performance and clustering accuracy. As the anomaly proportion increases, the AUC gradually degrades, with a drastic drop beyond 50%, where the dominant cluster assumption no longer holds. This is further evidenced by a significant decrease in the clustering accuracy.
Figure 4: Assessing the validity of anomaly scarcity assumption.
Few-shot versus unsupervised paradigms.
The results presented in Table 3 compare pretrained visual encoders and source-only detectors with different adaptation paradigms, i.e., few-shot, unsupervised, and supervised (oracle). The source-only finetuned model improves slightly over the pretrained ResNet50 visual encoder [15] but still has lower performance than the adaptation approaches, achieving an average AUC of 62.84%. Among the few-shot methods, our few-shot variant outperforms BiOST [7] and MsRA [22], achieving the highest AUC of 71.93%, which is comparable to the performance of a pretrained CLIP-ViT-B32 visual encoder. However, our unsupervised adaptation method surpasses all these models, with an average AUC of 75.54% indicating its ability to effectively mitigate domain gaps without relying on labeled target data. This can be explained by the fact that after clustering, our model has access to more representative normal target data than few-shot models, hence better generalizing to the target normal class. On the other hand, the Oracle, which has access to the target labels, achieves the highest performance (81.65%). The small gap between our unsupervised method and the oracle demonstrates the effectiveness of our approach even without supervision.
Normal class Zero-shot DSVDD src fine- tuned only Few-shot Adaptation Unsup. Adapt. Ours Super- (Adap. vised)
R50 CLIP BiOST MsRA Ours
CAD $\rightarrow$ Real
Aero. 41.05 74.97 67.71 36.80 81.56 81.55 84.86 90.91
Bicycle 67.35 90.28 65.12 59.20 68.45 74.58 81.45 81.73
Bus 28.58 42.27 66.01 47.90 68.12 72.26 82.17 72.16
Car 32.48 64.16 78.65 53.80 69.44 82.78 62.76 68.42
Horse 68.81 75.48 67.24 58.00 68.77 80.17 83.52 88.70
Knife 67.78 95.28 62.43 54.10 70.39 71.52 68.82 78.90
Motor. 60.07 82.25 69.45 58.10 65.64 80.16 91.15 83.46
Person 71.69 56.26 42.11 58.70 59.18 51.24 69.68 85.19
Plant 62.47 89.65 57.77 42.10 65.81 71.46 70.58 82.63
Skate. 85.00 91.52 60.70 41.60 61.30 63.17 83.71 83.73
Train 30.13 57.74 54.75 52.40 69.73 60.62 69.98 85.11
Truck 26.05 45.08 62.08 43.10 59.05 73.67 57.84 78.91
Avg. 53.45 72.08 62.84 50.48 67.28 71.93 75.54 81.65
$\pm$±std $\pm$19.57 $\pm$17.83 $\pm$8.55 $\pm$7.55 $\pm$5.79 $\pm$9.08 $\pm$9.80 $\pm$6.11
Table 3: AUC (%) performance on the target domain of our UDA anomaly detector on VisDA [28] compared with various adaptation paradigms (from zero-shot, i.e., pretrained Visual encoders, few-shot, to supervised, i.e., Oracle).
w/ Adaptation w/ Clustering w/ CLIP $\psi$ AUC (%)
62.84$\pm$8.55
@@ 6.3 Additional Experiments
Ablation on the framework components.
Table 4 provides the results obtained when each component, namely the adaptation loss, the dominant cluster identification through clustering, the use of an auxiliary visual encoder $\psi(\mathcal{D}^{t})$ or the trainable features $f(\mathcal{D}^{t})$. The results show that without adaptation, a model trained only on source data generalizes poorly to the target domain with only 62.84%. Direct adaptation of the source and the unlabeled target without clustering leads to inconsistent results, indicating low generalization capabilities to the target domain. Introducing clustering results in a significant performance boost. This can be seen when clustering is applied to the original representations of the feature extractor, as the performance improves by +5.63%, highlighting the importance of identifying the dominant cluster prior to alignment. Note that our method still outperforms the best few-shot baseline MsRA [22] (68.47% vs 67.28%) with just clustering and alignment (i.e. w/o CLIP) across all the VisDA classes. Finally, the best results are achieved when all components are combined. This setup boosts the average AUC to 75.54% on all VisDA classes. The substantial performance gains can be attributed to CLIP’s rich visual features, which, together with clustering and alignment, help achieve a more robust anomaly detector capable of better handling domain shift.
Dataset w/o Adapt. (Src Only) w/ Adaptation
Kmeans GMM Meanshift kNN
VisDA 62.84$\pm$08.55 75.54$\pm$10.23 72.24$\pm$08.81 74.14$\pm$07.44 71.65$\pm$06.68
[…]
Table 5: Ablation on different clustering techniques. GMM and K-means use $10$ components for VisDA and $2$ for other datasets. k=2 for VisDA and k=1 for the remaining datasets.
Clustering methods.
We compare different clustering techniques on three UDA benchmarks in Table 5. The first observation we make is that any type of clustering improves the performance. K-means and GMM have comparable results, without one clearly and consistently outperforming the other across datasets and adaptation directions. Meanshift clustering offers a performance increase compared to source-only models. However, its performance remains lower than that of the other clustering methods. In our experiments, we chose K-means clustering as it achieves comparable performance to GMM while requiring fewer parameters and simpler optimization. We further investigate the optimal number of K-Means components, as shown in Figure 5. The figure indicates that using 8 to 10 components yields the highest performance, with an AUC of approximately 75-76%. Decreasing the number of components would gradually degrade the performance. This suggests that a lower number of clusters may not capture the characteristics of the majority class, leading to inaccurate clustering and thus negatively impacting the generalization of the anomaly detection model across domains.
Figure 5: K-Means[14] components variation on VisDA.
$f$: ResNet50 + $\psi$: CLIP-ViT-B32
DSVDD [33] MSC [31]
ZS Src UDA ZS Src UDA
53.45 62.84 75.54 74.39 72.87 81.10
(+12.7$\uparrow$) (+8.23$\uparrow$)
Table 6: Our UDA approach on two anomaly detection methods: DSVDD [33] and Mean-shifted Contrastive Learning (MSC) [31]. ZS and Src mean Zero-shot and Source only, respectively.
Beyond DSVDD.
To assess whether our alignment approach applies to other unsupervised AD methods, we replace DSVDD with Mean-shifted Contrastive loss (MSC) [31] in Table 6. MSC adapts contrastive loss to the one-class setting by shifting augmented representations of the normal samples toward the mean of pretrained normal features, preserving their compactness. Similar to [32, 6], it uses a kNN density estimator to detect anomalies. Our results suggest that both methods benefit from the adaptation, as a consistent average improvement of +12.7% and +8.23% is seen across all the 12 classes of VisDA.
Alignment strategies.
@@
Table 8 compares the results of GNL [4] with our proposed UDA method on the PACS dataset [20], with Photo as the source and Art, Cartoon, and Sketch as the target domains. It can be seen that our UDA approach consistently outperforms GNL [4], particularly on Cartoon and Sketch domains. This suggests that unlike DG methods, which aim to generalize to any unseen domain solely by training on the source domain, UDA can be more effective for semantic UAD since it exposes the model to the target domain during training, even if it is unlabeled.
Adapt. Type Method Ph. $\rightarrow$ Art Ph. $\rightarrow$ Cartoon Ph. $\rightarrow$ Sketch
None Source only 64.06 64.08 57.35
DG GNL [4] 65.62 67.96 62.39
DA MsRA [22] (Few-shot) 71.43 69.89 61.87
Ours (Unsup.) 67.20 75.35 74.04
Table 8: AUC (%) of Domain Generalization (DG) for anomaly detection, trained ONLY on the source domain Photo (Ph.) and tested on unseen domains. DA means Domain Adaptation.